The nitrous acid/sodium nitrite conjugate pair has been chosen to prepare a buffer solution. What should the concentration ratio of NO2-/HNO2 be if the desired pH of this buffer is 4.00? (Ka of HNO2 = 4.27 x 10-4)
Answer #1

use:

pKa = -log Ka

= -log (4.27*10^-4)

= 3.37

use:

pH = pKa + log {[conjugate base]/[acid]}

4 = 3.37+log {[NO2-]/[HNO2]}

log {[NO2-]/[HNO2]} = 0.6304

[NO2-]/[HNO2] = 4.27

Answer: 4.27